MAT 145: Topics In Contemporary Math
QUESTION 11

What is 7 * 5 on a 12-hour clock?


Found 2 solutions by ikleyn, math_tutor2020:
Answer by ikleyn(52813)

.

7*5 = 11  (mod 12).    ANSWER

Solved, answered and explained.



Answer by math_tutor2020(3817)


Method 1: Quotient and remainder
7*5 = 35
35/12 = 2 remainder 11
Imagine you had 35 cookies and 12 friends.
Each friend gets 2 whole cookies (quotient), meaning they eat 2*12 = 24 cookies so far.
That leaves 35-24 = 11 cookies left over (remainder).
We're one cookie short of everyone able to get 3 whole cookies.

Method 2: Repeated subtraction
7*5 = 35
35-12 = 23
23-12 = 11
Recall that the process of long division involves repeated subtraction.
Notice there are two lines of subtracting by 12, to indicate the quotient of 2 mentioned earlier.

Method 3: Modular arithmetic
7*2 = 14 = 2 (mod 12)
7*4 = 2*(7*2) = 2*2 = 4 (mod 12)
7*5 = 7+7*4 = 7+4 = 11 (mod 12)
Ignore method 3 if your teacher has not covered this topic yet.
Even if your teacher introduced this notation style, it might still be a bit cryptic. So feel free to pick your favorite method.
Modular arithmetic is considered clock arithmetic because the numbers wrap around at some point. We only worry about the remainder.


Answer: 11